OpenStudy (anonymous):

Multiplying Polynomials by Monomials 3x (2x – 1) + x (3x – 7)

OpenStudy (anonymous):

any ideas for this one? you have to use the distributive property for both terms

OpenStudy (anonymous):

\[3x(2x-1)+x(3x-7)\] \[=3x\times 2x-3x\times 1+x\times 3x-7\times x\] or \[6x^2-3x+3x^2-7x\] then combine like terms

OpenStudy (anonymous):

do the exponents stay thee same